# On Modular Invariants of Truncated Polynomial Rings in low ranks

### v1

@@ Abstract.
We verify the conjectures due to Lewis, Reiner and Stanton about the Hilbert series of the invariant ring of the truncated polynomial ring for all parabolic subgroups up to rank $3$. This is done by constructing an explicit set of generators for each invariant ring in question. We also propose a conjecture concerning the action of the Steenrod algebra and the Dickson algebra on a certain naturally occurring filtration of the invariant ring under the general linear group.
@@
Truncated Polynomial Ring, Modular invariant, Dickson invariants, Steenrod algebra, q-binomial, q-multinomial, finite general linear group, Frobenius power
@@ 1. Introduction
Let $\mathbb{F}_{q}$ denote the finite field of $q$ elements, where $q$ is a power of a fixed prime number $p$. Let $I$ denote the ideal generated by the $m$th Frobenius power $x_{1}^{q^{m}},x_{2}^{q^{m}},\ldots,x_{n}^{q^{m}}$, of the polynomial algebra $S=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$. We denote by ${\mathcal{Q}}_{m}(n)$ or simply ${\mathcal{Q}}_{m}$, if $n$ is fixed, the quotient ring $S/I$. The general linear group $G=\operatorname{GL}_{n}(\mathbb{F}_{q})$ acts on $S$ by linear substitutions of variables. Since $I$ is a $G$-invariant ideal, ${\mathcal{Q}}_{m}(n)$ inherits a $G$-action from $S$. This paper discusses the problem of computing invariants of the truncated ring ${\mathcal{Q}}_{m}(n)$ under the action of various parabolic subgroups of $G$.
This problem arises in various contexts, in connection with questions in algebraic combinatorics ([14], [15], [10], [4]), algebraic topology ([1], [7], [9]) and number theory [2]. In particular, taking the limit $m$ to infinity yields results about the cofixed space $S_{G}$ of the polynomial ring under the action of the general linear group. This problem is surprisingly still open in modular invariant theory.
Let $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ be a composition of $n$ and denote by $P(\alpha)$ the corresponding parabolic subgroup. In [10], Lewis, Reiner, and Stanton proposed a marvelous series of conjectures about the Hilbert-Poincaré series of the invariant ring ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ for arbitrary $m$ and any composition $\alpha$ of $n$ in terms of a new combinatorial object called the $(q,t)$-binomial coefficient introduced in [15]:
@@
([10, Conjecture 1.5]) The Hilbert-Poincaré series of the ring of invariants of ${\mathcal{Q}}_{m}(n)$ under the action of the parabolic subgroup $P(\alpha)$ is:
@@
In particular, when $\alpha=(n)$, they conjectured that
@@
([10, Conjecture 1.2]) The Hilbert-Poincaré series of the ring of invariants of ${\mathcal{Q}}_{m}(n)$ under the action of the general linear group $G=\operatorname{GL}_{n}(\mathbb{F}_{q})$ is:
@@
These conjectures are inspired by properties of the $q$-Catalan and $q$-Fuss Catalan numbers, connecting Hilbert-Poincaré series for certain invariant subspaces with the representation theory of rational Cherednik algebras for Coxeter and complex reflection groups. Some partial results about these two conjectures have been obtained (see Drescher-Shepler [4], Goyal [5] and the original paper [10]). Still, how these combinatorial objects occur in the series remains quite mysterious.
In our previous work [6], we have proved their parabolic conjecture for the minimal parabolic subgroup by constructing an explicit basis. We also refined their conjectures by proposing an explicit $\mathbb{F}_{q}$-basis for ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ as follows:
@@
A basis for ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ is given inductively by the following formula:
$\mathcal{B}_{m}(\alpha)=\{\delta_{s+1;m}^{\alpha_{1}-s}(f\cdot\Phi^{s}g)\colon 0\leq s\leq\min(\alpha_{1},m),f\in\Delta^{m}_{s},g\in\mathcal{B}_{m-s}(\alpha_{2},\ldots,\alpha_{\ell})\}.$
Here $\Delta^{m}_{s}$ is a certain subspace of the Dickson algebra ([3]) of all $\operatorname{GL}_{s}$-invariant polynomials and $\delta_{a;b}$ is the operator constructed in [6]. The Frobenius-like operator $\Phi$ is given by the rules:
  1. (i)
    $\Phi(\delta_{a;b})=\delta_{a+1;b+1}$;
  2. (ii)
    $\Phi(Q_{r,i})=Q_{r+1,i+1}$ and expand multiplicatively.
The reader of [6] may notice that in this conjecture, we have used the core subspace $\Delta^{m}_{s}$ of the Dickson algebra instead of $\nabla^{m}_{s}$ in the original conjecture. The two candidates can be used interchangeably, but $\Delta^{m}_{s}$ is more convenient for explicit calculation, as we will see later.
In the case $\alpha=(n)$ so that $P(\alpha)=G$ is the full general linear group, we have
@@
forms a basis for the $\mathbb{F}_{q}$-vector space ${\mathcal{Q}}_{m}^{G}(n)$.
It is a simple counting problem that our proposed bases have the correct series as predicted. Hence our conjectures refine the two conjectures proposed by Lewis, Reiner, and Stanton. Furthermore, they also explain where the summands of the Hilbert-Poincaré series occur in the original conjectures: The summands of $C_{n,m}(t)$ are just the Hilbert-Poincaré series for $\Delta^{m}_{s}$, $0\leq s\leq\min(m,n)$, shifted by an appropriate degree because of the iterated delta operator.
In [6], we have verified the refined parabolic conjecture 1.3 for the Borel subgroup, that is when $\alpha=1^{n}$. The aim of this note is to verify our conjectures for other parabolic subgroups, including the full general linear group, for $n\leq 3$. As an immediate corollary, the parabolic conjecture of Lewis, Reiner and Stanton is true in these cases.
@@
In this paper, we choose to work as explicitly as possible to demonstrate the underlying mechanism of our construction and the additional structure of our basis. The computation reveals an interesting property of the invariant rings considered as a module over the Dickson algebra. We also describe the action of the Steenrod algebra. In subsequent work, we hope to return to this problem in greater generality.
When $n=2$, the only parabolic subgroup other than the Borel subgroup is $\operatorname{GL}_{2}$ itself, and the parabolic conjecture has already been verified in the original paper [10]. However, the authors worked with the dual situation for cofixed spaces and used a particular property of the bivariate case to avoid the truncated ring. Thus, it is unclear whether the invariant subspace ${\mathcal{Q}}_{m}^{G}$, even in rank $2$, can be described explicitly using their method.
In [5, Proposition 4.1], Goyal constructed a family of $G$-invariant polynomials in ${\mathcal{Q}}_{m}(2)$, which he denoted as $y_{k^{\prime}}$ ($0\leq k^{\prime}\leq\frac{q^{m}-q}{q-1}$). It turns out that this family coincides with the $G$-invariant family $\delta_{2}(\Delta^{m}_{1})$ in our notation.
To simplify the notation, we will often drop the index $n$ in ${\mathcal{Q}}_{m}(n)$ when the number of variables is well-understood from the context. Moreover, since the Frobenius power $m$ is fixed, $\delta_{s}$ will be always used in place of $\delta_{s;m}$ throughout the paper. We will also write $(i_{1},i_{2})\in\Delta^{m}_{2}$ instead of $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}\in\Delta^{m}_{2}$. Finally we denote by $[a]_{q}$ the $q$-integer $(q^{a}-1)/(q-1)$.
@@
The set $\mathcal{B}_{m}(2)$ consisting of 3 families below forms a basis for the invariant subspace ${\mathcal{Q}}_{m}^{G}(2)$ of the full general linear group.
@@
In rank 3, other than the Borel subgroups, there are 2 proper parabolic subgroups. Their corresponding rings of invariants are described in the next 2 propositions.
@@
Since the case of the Borel subgroup, which corresponds to the composition $(1,1,1)$ of $n=3$, has been discussed in full generality in [6], the Propositions above combined imply the main Theorem 1.5.
Note that when $m$ is small compared to $n$, all results remain valid with proper interpretation. For example if $n=3$ and $m=2$, then $\Delta^{2}_{3}$ is empty, and $\delta_{3;2}(\Delta^{2}_{2})$ just consists of a single element, which is $1$ in degree 00. To avoid too many special cases (whose proof are not complicated), we will generally assume that $m$ is sufficiently large, comparing to $n$.
In the course of proving these results, we have also computed the structure of these invariant rings, considered as a module over the Dickson algebra, ${\mathcal{D}}_{n}$, and a module over the Steenrod algebra $\mathcal{A}$. At least in the case we have computed, our basis yields a remarkable filtration, namely the subspaces $\mathcal{F}_{n,k}$, for $0\leq k\leq n$, of ${\mathcal{Q}}_{m}^{G}(n)$ spanned by $\delta_{s+1}^{n-s}(\Delta^{m}_{s})$ for all $s\leq k$. We have the following conjecture about this filtration:
@@
For each $1\leq k\leq\min(m,n)$, the subspace $\mathcal{F}_{n,k}$ given by
$\mathcal{F}_{n,k}=\operatorname{Span}\{\delta_{s+1}^{n-s}(f)\colon f\in\Delta^{m}_{s},0\leq s\leq k\},$
is an $\mathcal{A}$-submodule as well as a ${\mathcal{D}}_{n}$-submodule of ${\mathcal{Q}}_{m}^{G}(n)$. Moreover, $\mathcal{F}_{n,k}$ is annihilated by the Dickson invariants $Q_{n,0},Q_{n,1}\ldots Q_{n,n-k-1}$.
This conjecture suggests that the filtration $\mathcal{F}_{n,*}$ may have another description of topological nature. We can prove the following:
@@
The layout of this paper is as follows: Section 2 recalls basic results about modular invariant theory and the key construction from our previous work.
In Section 3, we give a lower bound on the total dimension of the invariant subspaces. The following four Sections are devoted to demonstrating the four Propositions from 1.6 to 1.9 above. The last section discusses the action of the Dickson algebra and the Steenrod algebra and the proof of Theorem 1.11.
@@ Acknowledgement
This work was partially completed while the second and third authors visited the Vietnam Institute for Advanced Study in Mathematics (VIASM). We want to thank Vo Thanh Tung for some computer calculations in the initial stage of this project.
@@ 2.1. Upper triangular invariants and Dickson invariants
The canonical projection $S\to{\mathcal{Q}}_{m}$ is clearly a $\operatorname{GL}$-map, hence for each composition $\alpha$ of $n$, there is an induced $\mathbb{F}_{q}$-algebra homomorphism
$S^{G}\subset S^{P(\alpha)}\to{\mathcal{Q}}_{m}^{P(\alpha)},$
making the invariant ring ${\mathcal{Q}}_{m}^{P(\alpha)}$ into a module over $S^{G}$. Thanks to the work of Dickson [3] since 1911, the structure of the algebra $S^{G}$, which we now call the Dickson algebra, is well-known, and plays a significant role in our investigation.
@@
It is well-known that the invariant ring $S^{B}$ under the Borel subgroup of upper triangular matrices is a polynomial algebra generated by $V_{i}^{q-1}$, $1\leq i\leq n$. Let $X$ be an indeterminate and recall the fundamental equation
@@
The polynomials $Q_{n,i}=Q_{n,i}(x_{1},\ldots,x_{n})$ are evidently $\operatorname{GL}_{n}$-invariant. They are called the Dickson invariants, because of the following celebrated theorem due to Dickson:
@@
The Dickson invariants $Q_{n,i}$ can be described in a more explicit form as follows. For any $n$-tuple of non-negative integers $(r_{1},\dots,r_{n})$, put
$[r_{1},\dots,r_{n}]:=\det(x_{i}^{q^{r_{j}}})_{1\leq i,j\leq n}$
and let $L_{n}=L_{n}(x_{1},\ldots,x_{n})$ denote $[0,1\ldots,n-1]$. We then have for $1\leq i\leq n$
$Q_{n,i}=[0,\ldots,\hat{i},\ldots,n]/L_{n},$
In particular, $Q_{n,n}=1$ and $Q_{n,0}=L_{n}^{q-1}$. By convention, we set $Q_{n,i}=0$ if $i<0$ or $i>n$. The following relations are well-known (see for example [13]):
$L_{n}=V_{1}\cdots V_{n},\quad Q_{n,i}=Q_{n-1,i-1}^{q}+Q_{n-1,i}V_{n}^{q-1}.$

### v2

# On Modular Invariants of the Truncated Polynomial Rings in low ranks

@@ Abstract.
We verify the conjectures due to Lewis, Reiner, and Stanton about the Hilbert series of the invariant ring of the truncated polynomial ring for all parabolic subgroups up to rank $3$. This is done by constructing an explicit set of generators for each invariant ring in question. We also propose a conjecture concerning the action of the Steenrod algebra and the Dickson algebra on a certain naturally occurring filtration of the invariant ring under the action of the general linear group.
@@
Truncated Polynomial Ring, Modular invariants, Dickson invariants, Steenrod algebra, q-binomial, q-multinomial, finite general linear group, Frobenius power
@@ 1. Introduction
Let $\mathbb{F}_{q}$ denote the finite field of $q$ elements, where $q$ is a power of a fixed prime number $p$. Let $I_{m}(n)=(x_{1}^{q^{m}},\ldots,x_{n}^{q^{m}})$ denote the $m$-th Frobenius power ideal in the polynomial algebra $S(n)=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$. Denote by ${\mathcal{Q}}_{m}(n)$ the truncated polynomial ring $S(n)/I_{m}(n)$. The general linear group $G_{n}=GL_{n}(\mathbb{F}_{q})$ acts on $S(n)$ by linear substitutions of variables. Since $I_{m}(n)$ is a $G_{n}$-invariant ideal, ${\mathcal{Q}}_{m}(n)$ inherits a $G_{n}$-action from $S(n)$. As we will be working with a fixed number of generators $n$, we will usually omit $n$ from the notation.
This paper addresses the problem of computing invariants of the truncated ring ${\mathcal{Q}}_{m}(n)$ under the action of various parabolic subgroups of $G_{n}$. This problem arises in several contexts, including algebraic combinatorics ([11], [12], [9], [4]), algebraic topology ([1], [7], [8]) and number theory [2]. In particular, as shown in [9, Conjecture 1.3], knowing the Hilbert series of ${\mathcal{Q}}_{m}(n)^{G}$ provides the Hilbert series of the cofixed space $S_{G}$ of the polynomial ring under the action of the general linear group. Surprisingly, while the invariant ring $S^{G}$ has been completely determined since the beginning of the 20th century, by the fundamental work of Dickson [3], the problem of determining $S_{G}$ is still an open problem in modular invariant theory.
For a positive integer $n$, a composition of $n$ of length $\ell$ is an ordered set $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ of positive integers $\alpha_{i}$ such that $|\alpha|=\alpha_{1}+\ldots+\alpha_{\ell}=n$. We will also consider weak compositions, which are ordered sets of non-negative integers. The set of weak compositions of a given length is partially ordered by declaring that $\beta\leq\alpha$ iff $\beta_{i}\leq\alpha_{i}$ for all $i$. If $\alpha$ is a composition of $n$, we denote by $P(\alpha)$ the corresponding parabolic subgroup of $G_{n}$. In [9], Lewis, Reiner, and Stanton proposed a marvelous set of conjectures about the Hilbert series of the invariant ring ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$ for arbitrary $m$ and any composition $\alpha$ of $n$, expressed in terms of a new combinatorial object called the $(q,t)$-binomial coefficient, introduced in [12]:
@@
([9, Parabolic Conjecture 1.5]) The Hilbert series of the ring of invariants of ${\mathcal{Q}}_{m}(n)$ under the action of the parabolic subgroup $P(\alpha)$ is:
@@
In particular, in the most interesting case where $\alpha=(n)$, they conjectured that
@@
([9, Conjecture 1.2]) The Hilbert series of the ring of invariants of ${\mathcal{Q}}_{m}(n)$ under the action of the general linear group $G_{n}$ is:
@@
These conjectures are inspired by properties of the $q$-Catalan and $q$-Fuss-Catalan numbers, connecting Hilbert series of certain invariant subspaces with the representation theory of rational Cherednik algebras for Coxeter and complex reflection groups. Some partial results about these two conjectures have been obtained (see Drescher-Shepler [4], Goyal [5], and the original paper [9]). Nevertheless, the role of these combinatorial objects in the series remains quite mysterious.
In our previous work [6], we proved their parabolic conjecture for the minimal parabolic subgroup $B$ by constructing an explicit basis for the $\mathbb{F}_{q}$-vector space ${\mathcal{Q}}_{m}(n)^{B}$ for all $m$ and $n$. We also refined their conjectures by proposing an explicit $\mathbb{F}_{q}$-basis for ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$.
@@
A basis for the space of $P(\alpha)$-invariants of ${\mathcal{Q}}_{m}(n)$ is given by the set $\mathcal{B}_{m}(\alpha)$ consisting of elements of the form
$\delta_{B_{1}+1;m}^{\alpha_{1}-\beta_{1}}\bigg{(}f_{1}\delta_{B_{2}+1;m}^{\alpha_{2}-\beta_{2}}\big{(}\cdots f_{\ell-1}\delta_{B_{\ell}+1;m}^{\alpha_{\ell}-\beta_{\ell}}(f_{\ell})\big{)}\bigg{)}$
where $\beta\leq\alpha,|\beta|\leq m$, $B_{i}=\beta_{1}+\ldots+\beta_{i}$ (by convention, $B_{0}=0$) and $f_{i}\in\Phi^{B_{i-1}}\Delta^{m-B_{i-1}}_{\beta_{i}}\subseteq\Delta^{m-B_{i}}_{B_{i}}$.
Here $\delta_{a;b}$ is the operator constructed in [6] and $\Delta^{m}_{s}$ is a certain subspace of the Dickson algebra $\mathcal{D}_{s}$ ([3]) of all $GL_{s}$-invariant polynomials:
$\Delta^{m}_{s}\subset\mathcal{D}_{s}=\mathbb{F}_{q}[Q_{s,s-1},\ldots,Q_{s,1},Q_{s,0}].$
The Dickson invariants $Q_{s,i}$ are crucial to our investigation and will be reviewed in detail in Section 2. The Frobenius-like operator $\Phi$ is a ring map from $\mathcal{D}_{s}$ to $\mathcal{D}_{s+1}$ which sends $Q_{s,i}$ to $Q_{s+1,i+1}$.
The reader of [6] may notice that in this conjecture, we describe a member of the conjectural generating set $\mathcal{B}_{m}(\alpha)$ explicitly rather than by induction. In addition, we make use of the core subspace $\Delta^{m}_{s}$ (see 2.2 below) of the Dickson algebra instead of $\nabla^{m}_{s}$ in the original conjecture. The two candidates for core spaces can be used interchangeably, but $\Delta^{m}_{s}$ is more convenient for explicit calculation, as we will see later.
In the case $\alpha=(n)$ so that $P(\alpha)$ is the full general linear group, we have
@@
forms a basis for the $\mathbb{F}_{q}$-vector space of $G_{n}$-invariants of ${\mathcal{Q}}_{m}(n)$.
It is a simple counting problem that our proposed bases have the correct Hilbert series as predicted. Hence our conjectures refine the two conjectures proposed by Lewis, Reiner, and Stanton. Furthermore, they also explain where the summands of the Hilbert series occur in the original conjectures: The summands of $C_{n,m}(t)$ are just the Hilbert series of $\Delta^{m}_{s}$, $0\leq s\leq\min(m,n)$, shifted by an appropriate degree due to the iterated delta operator.
In [6], we have verified the refined parabolic conjecture 1.3 for the Borel subgroup, that is when $\alpha=1^{n}$. The aim of this paper is to verify our conjectures for other parabolic subgroups, including the full general linear group, for $n\leq 3$. As an immediate corollary, the parabolic conjecture of Lewis, Reiner and Stanton is true in these cases.
@@
In this paper, we choose to work as explicitly as possible to illuminate the underlying mechanism of our construction and to highlight the additional structure present in our basis. Our computations reveal an interesting property of the invariant rings when considered as modules over the Dickson algebra. We also describe the action of the Steenrod algebra. In future work, we hope to revisit this problem in greater generality.
When $n=2$, the only parabolic subgroup other than the Borel subgroup is $G_{2}$ itself, and the parabolic conjecture has already been verified in the original paper [9]. However, the authors worked with the dual situation involving cofixed spaces and used a specific property of the bivariate case to avoid the truncated ring. As a result, it remains unclear whether the invariant subspace of ${\mathcal{Q}}_{m}(n)$, even in rank $n=2$, can be described explicitly using their method.
In [5, Corollary 4.3], Goyal constructed a family of $G_{n}$-invariant polynomials in ${\mathcal{Q}}_{m}(n)$ which corresponds to the family $\delta_{2}^{n-1}(\Delta^{m}_{1})$ in our notation.
To simplify the notation, we will often drop the index $n$ in ${\mathcal{Q}}_{m}(n),S(n)$, and $G_{n}$ when the number of variables is well-understood from the context. If the Frobenius power $m$ is fixed, we write $\delta_{s}$ instead of $\delta_{s;m}$. It is also convenient to write $(e_{1},\ldots,e_{s})\in\Delta^{m}_{s}$ to indicate a Dickson monomial $Q_{s,s-1}^{e_{1}}\ldots Q_{s,0}^{e_{s}}$ in $\Delta^{m}_{s}$. For any nonnegative integer $a$, we denote by $[a]_{q}$ the $q$-integer $(q^{a}-1)/(q-1)$ with the convention that $[0]_{q}=1$.
@@
The set $\mathcal{B}_{m}(2)$ consisting of 3 families below forms a basis for the invariant subspace ${\mathcal{Q}}_{m}(2)^{G}$ of the full general linear group.
@@
In rank 3, other than the Borel subgroups, there are two proper parabolic subgroups. Their corresponding rings of invariants are described in the next two propositions.
@@
Since the case of the Borel subgroup, which corresponds to the composition $(1,1,1)$ of $n=3$, was treated in [6], the propositions stated above together imply the main Theorem 1.5.
It is worth noting that when $m$ is small relative to $n$, all results still hold, provided they are interpreted appropriately. For example, when $n=3$ and $m=2$, the set $\Delta^{2}_{3}$ is empty, and $\delta_{3;2}(\Delta^{2}_{2})$ consists of just a single element, namely $\delta_{3;2}(1)=1\in\mathbb{F}_{q}$ in degree 00. To avoid excessive attention to these special cases (whose proofs are straightforward), we will generally assume that $m$ is sufficiently large compared to $n$ (in fact, $m\geq n$ is enough!).
In the process of establishing these results, we have also computed the structure of these invariant spaces considered as modules over the Dickson algebra and the mod $q$ Steenrod algebra.
Definition 1.10.
For each $0\leq k\leq\min(m,n)$, let $\mathcal{F}_{n,k}$ denote the subspace of ${\mathcal{Q}}_{m}(n)^{G}$ spanned by the subsets $\delta_{s+1}^{n-s}(\Delta^{m}_{s})$ where $0\leq s\leq k$:
$\mathcal{F}_{n,k}=\operatorname{Span}\{\delta_{s+1}^{n-s}(f)\colon f\in\Delta^{m}_{s},0\leq s\leq k\}.$
These subspaces form an increasing filtration of ${\mathcal{Q}}_{m}(n)^{G}$, starting with the one-dimensional subspace $\mathcal{F}_{n,0}$, which is spanned by the top-degree class $(x_{1}\ldots x_{n})^{q^{m}-1}$. At least in the cases we have computed, this filtration is exhaustive and exhibits remarkable properties. We propose the following:
@@
For each $1\leq k<\min(m,n)$, $\mathcal{F}_{n,k}$ is an $\mathcal{A}$-submodule as well as a ${\mathcal{D}}_{n}$-submodule of ${\mathcal{Q}}_{m}(n)^{G}$. Moreover, $\mathcal{F}_{n,k}$ is annihilated by the Dickson invariants $Q_{n,0},Q_{n,1}\ldots Q_{n,n-k-1}$.
This conjecture suggests that the filtration $\mathcal{F}_{n,*}$ may admit an alternative description of a topological nature. We are able to verify this conjecture up to rank $3$:
@@
The layout of this paper is as follows: Section 2 recalls basic results in modular invariant theory and the key constructions from our previous work. In Section 3, we establish a lower bound on the total dimension of the invariant subspaces. The subsequent four sections are devoted to proving Propositions 1.6 through 1.9. The final section discusses the actions of the Dickson algebra and the Steenrod algebra, and presents the proof of Theorem 1.12.
@@ Acknowledgement
This work was partially completed while the second and third authors were visiting the Vietnam Institute for Advanced Study in Mathematics (VIASM). The third author gratefully acknowledges support from Hung Vuong University’s Fundamental Research Program (Project No. 03/2024). We would like to thank Vo Thanh Tung for providing some computer calculations during the initial stages of this project. We are also grateful to the anonymous referee for their careful reading and numerous suggestions, which significantly improved the exposition.
@@ 2.1. Upper triangular invariants and Dickson invariants
The canonical projection $S\to{\mathcal{Q}}_{m}$ is clearly a $G$-equivariant map. Thus, for each parabolic subgroup of $G$, there is an induced $\mathbb{F}_{q}$-algebra homomorphism
$S^{G}\subset S^{P}\to{\mathcal{Q}}_{m}^{P},$
which makes the invariant ring ${\mathcal{Q}}_{m}^{P}$ into a module over $S^{G}$. The structure of the invariant ring $S^{G}$, which is the Dickson algebra [3], is well-known and plays a significant role in our investigation.
@@
It is well-known that the invariant ring $S^{B}$ under the Borel subgroup of upper triangular matrices is a polynomial algebra generated by $V_{1}^{q-1},V_{2}^{q-1},\ldots,V_{n}^{q-1}$ (see for example, [10, Theorem 3.4]). Let $X$ be an indeterminate, and recall the fundamental equation
@@
The polynomials $Q_{n,i}=Q_{n,i}(x_{1},\ldots,x_{n})$ are evidently $G$-invariant, and are called Dickson invariants because of the following celebrated theorem due to Dickson:
@@
The Dickson invariants $Q_{n,i}$ can be described more explicitly as follows. For any $n$-tuple of non-negative integers $(r_{1},\dots,r_{n})$, put
$[r_{1},\dots,r_{n}]=\det(x_{i}^{q^{r_{j}}})_{1\leq i,j\leq n}.$
In particular, let $L_{n}=L_{n}(x_{1},\ldots,x_{n})$ denote $[0,1\ldots,n-1]$. We then have
$Q_{n,i}=[0,\ldots,\hat{i},\ldots,n]/L_{n}.$
In particular, $Q_{n,n}=1$ and $Q_{n,0}=L_{n}^{q-1}$. By convention, we set $Q_{n,i}=0$ if $i<0$ or $i>n$. The following relations are well-known (see for example, [15, Proposition 1.3]):
Proposition 2.1.
Let $n$ be a positive integer and $0\leq i\leq n$. The following relations hold:
  • (i)
    $L_{n}=V_{1}V_{2}\cdots V_{n}$.
  • (ii)
    $Q_{n,i}=V_{n}^{q-1}Q_{n-1,i}+Q_{n-1,i-1}^{q}$.